My cats' scratching post is 860mm high with a diameter of 80mm. What length of sisal rope will I need to cover the surface area of the post, if the diameter of the rope is 8mm
Answer by Alan3354(69443)

My cats' scratching post is 860mm high with a diameter of 80mm. What length of sisal rope will I need to cover the surface area of the post, if the diameter of the rope is 8mm
----------------------
The # of turns = 860/8 =~ 108
Each turn = 84*pi mm
--> 108*84*pi =~ 28500 mm
= 28.5 meters